13. An automobile accelerates uniformly at 8 m/s2. If the force causing the acceleration is 6000 N, what is the mass of the automobile?


14. Find the force required to increase the velocity of a 1.2 kg mass from 3 m/s to 8 m/s over a period of 4 seconds.


1
Expert's answer
2021-12-06T09:42:55-0500

13.

"m=\frac{F}{a}=\rm\frac{6000\: N}{8\: m/s^2}=750\: kg"

14.

"F=m\frac{\Delta v}{\Delta t}=\rm 1.2\: kg*\frac{5\: m/s}{4\: s}=1.5\: N"